OpenStudy (chihiroasleaf):

If \(x+ \frac{1}{y}=8\) and \(xy+ \frac{1}{xy}=38\), then the value of \(y+ \frac{1}{x}=....\)

OpenStudy (amistre64):

x + 1/y = 8 xy + 1/xy = 38 substitution might help yx + 1 = 8y xy = 8y - 1 therefore 8y - 1 + 1/xy = 38 8y^2 - y + 1/x = 38y 8y^2 - y-y +y+ 1/x = 38y y+ 1/x = 38y +2y-8y^2

OpenStudy (rational):

was just wondering if dot products might help a = (x, 1/y) b = (1, 1) c = (y, 1/x) a.b = 8 a.c = 38 b.c = ?

OpenStudy (amistre64):

that looks plausible as well :)

OpenStudy (rational):

not so sure, im stuck actually haha..

OpenStudy (amistre64):

its a good thought tho, id never of considered it thats for sure

Parth (parthkohli):

I have my own stupid trick.\[\dfrac{xy + 1}{y } = 8\tag{1}\]\[\dfrac{xy + 1}{x} = ~?\tag{2}\]\[\dfrac{(xy)^2 + 1}{xy} = 38\tag{3}\]Multiply 1 and 2:\[\dfrac{(xy)^2 + 1 + 2xy}{xy} = (3) + 2\]

OpenStudy (amistre64):

latex isnt coding for me so thats just an eyesore :P

OpenStudy (chihiroasleaf):

This is what I've got so far: \[xy+\frac{1}{xy}=38\] \[(xy)^2+1=38xy\] \[(xy)^2-38xy +1 =0\] \[xy =\frac{ 38 \pm \sqrt{1444-4}}{2} = \frac{ 38 \pm \sqrt{1440}}{2} = 19 \pm 6\sqrt{10}\]

OpenStudy (rational):

http://gyazo.com/08afd4459b71491ed74a314e087f263a

OpenStudy (chihiroasleaf):

@ParthKohli that's great.., how do you get the idea?

Parth (parthkohli):

OS is acting up. Anyway,\[(3) + 2 = 40 = 8 \times ?\]\[\Rightarrow ~? = 5\]

Parth (parthkohli):

Oh, I don't know either. I've noticed that my brain is coming up with ideas it couldn't earlier.

OpenStudy (adi3):

di u get it

mathslover (mathslover):

Well done @ParthKohli !